OpenStudy (anonymous):

Solve the equation. A. B. –12 C. D. –35

OpenStudy (anonymous):

Solve the equation. -5= c/7 A.) - 7/5 b.) -12 c.) -5/7 d.) -35 this is the full question.

OpenStudy (anonymous):

@robtobey can you help me with this one?

OpenStudy (anonymous):

-5 = c/7 Multiply each side by 7. 7(-5) = (c/7)7 -35 = c

OpenStudy (anonymous):

thank you :)